Exercise 22-6 The Palmer Acres Inn is trying to determine its break-even point during its off-peak season. The inn has 50 rooms that it rents at $70 a night. Operating costs are as follows: Salaries Utilities Depreciation Maintenance Maid service Other costs $5,800 per month $1,400 per month $1,100 per month $3,740 per montlh $14 per room $28 per room Determine the inns break-even point in number of rented rooms per month Break-even point rooms LINK TO TEXT LINK TO TEXT Determine the inns break-even point in dollars Break-even point Click if you would like to Show Work for this question:
Answer #1

Fixed cost = 5800+1400+1100+3740 = 12040

Break even point = Fixed cost/(Selling price per unit-variable cost per unit)

= 12040/(70-42)

Break even point = 430 rooms

Break even point = 430*70 = $30100